OpenStudy (anonymous):

40n^2+9n Please convert this to the "sum" formula.

OpenStudy (anonymous):

Any one?

OpenStudy (anonymous):

n(40n+9)

myininaya (myininaya):

a sum formula?

OpenStudy (anonymous):

Perhaps he means that is a "term" and he wants to "sum" the terms?

jhonyy9 (jhonyy9):

how you have thought it ? - like a*2 +2ab +b*2 = (a+b)*2 - ???

OpenStudy (anonymous):

I meant like a "sigma" sum but I couldn't see any way of doing it with the squared term there.

OpenStudy (anonymous):

Assuming an arithmetic series of n terms with common difference d and first term a, then the sum is equal to \[\frac{d}{2}n^2+\left(a+\frac{d}{2}\right)n\] Compare this expression to your answer. Once you have solved for d and a, then the sum will be in the form \[\sum_{i=1}^n(a+id)\]

OpenStudy (anonymous):

Well, yes except that its 40n^2 + 9n....

OpenStudy (anonymous):

\[\frac{d}{2}=40\] \[\left(a+\frac{d}{2}\right)=9\] Where is the problem?

OpenStudy (anonymous):

Set n=1, 2, 3... 49,169,387.... What is the sum Sn?

OpenStudy (anonymous):

The original question can be paraphrased as, if I'm not mistaken, "what is the sum that gives the equality \[\sum_{i=1}^nt_i = 40n^2+9n\] not "what is \[\sum_{i=1}^n(40i^2+9i)\]

OpenStudy (anonymous):

Well, that's what the original replies were about, what DID he mean and he hasn't returned to tell us so we all just guessing what he might have meant.